Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30
If x represents the number of quarters in the jar, which equation can be used to find the number of quarters?

Answer by ankor@dixie-net.com(22740)

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30
If x represents the number of quarters in the jar, which equation can be used to find the number of quarters?
:
(n = no. of nickels)
x + n = 150
n = (150 - x)
and
.25x + .05n = 28.30
.05n = (28.30 - .25x)
Multiply eq by 20
n = (566 - 5x)
:
find x using this equation
150 - x = 566 - 5x
-x + 5x = 566 - 150
4x = 416
x = 104 quarters